• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 10:40:26 on June 27, 2010

    Entry number 328725

    keyword=Experiment status

    Beam is back, we checked the position and the rates and are back to taking optics data.

    We are planning to do the pass change and polarity change around 5 PM. If think we will change the Gas Cerenkov PMT around that time or ask for the access when Optics plan is complete.

    We are still waiting to hear from Steve Suhring about the permission to go to higher beam hopefully we will get it during the day.



    A copy of this log entry has been emailed to: apex@jlab.org,kees@jlab.org